OpenStudy (anonymous):

S= (rL-a)/(r-1) solve for r

OpenStudy (anonymous):

S=(rL-a)/(r-1) Multliply both sides by (r-1) to get S(r-1) = rL-a Distribute Sr - S = rL - a Subtract rL and Add S to get Sr - rL = S - a Factor out r r(S - L) = S - a Divide out r = (S-a) / (S-L)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Your welcome! :)